OpenStudy (anonymous):

Find the area of the region bounded by the curves y=x^2 y=x^3 how do i find the intersection points?

OpenStudy (anonymous):

Set them equal:\[x^3 = x^2 \implies x^3 - x^2 = 0 \implies x^2(x - 1) = 0 \implies x_1 = 0, x_2 = 0, x_3 = 1\]

OpenStudy (anonymous):

so those are the x coordinates right?

OpenStudy (anonymous):

I got all that and the (0,0) but i dont get that the other point is (1,1)

OpenStudy (hunus):